OpenStudy (anonymous):

what is this um The length of a rectangle is 3 more than 4 times its width. The perimeter of the rectangle is 86 inches. What is the length of the rectangle? 23 in. 27 in. 31 in. 35 in.

OpenStudy (anonymous):

 Sorry if it's abit long, it's pretty simple though \[ b = 4 \cdot a + 3 \\ \text{perimeter} = 2\cdot b + 2 \cdot a \\ \text {perimeter} = 2 \cdot (4a + 3) + 2a \\ \text {perimeter} = 8a + 6 + 2a \\ \text {perimeter} = 10a + 6 \\ \text {perimeter} = 86_{inches}\\ 10a + 6 = 86 \\ 10 a = 80 \\ a = 8_{inches} \\ b = 4\cdot a + 3 = 4\cdot8 + 3 = 32 + 3 = 35_{inches} \]